THE MILLIMETER AND SUBMILLIMETER SPECTRUM OF BANH(X^{1}Σ^{+})

\underline{A. J. JANCZYK} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of BaNH in its X^{1}Σ^{+} ground electronic state has been recorded using millimeter/submillimeter direct absorption methods in the range 366 - 415 GHz, as well as that of its deuterium isotopomer. The molecules were produced in the presence of a d.c. discharge by the reaction of ammonia or ND_{3} and barium vapor, produced in the Broida - type oven. Transitions arising from the ground vibrational state and the excited vibrational bending mode (01^{1}0) were measured. The molecule appears to be linear, with B_{0}(BaNH) = 7984.549 MHz and B_{0}(BaND) = 7060.446 MHz. An r_{0} structure indicates r_{0}(BaN) = 2.078 Å and r_{0}(NH) = 1.011 Å.